OpenStudy (anonymous):

(HELP PLEASE) Given the equation of a line in standard form, Ax+By=C with B not equal to 0, which of the following represents the slope of this line? A. m=-B/A B. m= A/B C. m= C/B D. m= -A/B I do not think its C.

OpenStudy (dtan5457):

Ax is your slope. That means you have to move this to the other side so you get by=c-ax Whatever the value of y is, -ax must be divided by by.

OpenStudy (dtan5457):

by BY<<y

OpenStudy (dtan5457):

Example. Let's say it was 5y+10x=5 You would move the the x to the other side, right? 5y=5-10x To get your actual slope, 5y can't be 5y, it must be 5. So you divide 10x by 5. y=1-2x

OpenStudy (anonymous):

so it would be -A/B?

OpenStudy (dtan5457):

I believe so, yeah.

OpenStudy (anonymous):

Thank you.